Solve the following equations and check your results.5x + 9 = 5 + 3x

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

5x+9=5+3x
⇒5x-3x=5-9
⇒2x=-4
⇒x=-2
Putting the value of x in RHS and LHS we get,
5× (-2) + 9 = 5 + 3× (-2)
⇒-10+9=5+(-6)
⇒-1=-1
⇒LHS=RHS

Thanks